write an equation for
adult tickets to a play cost $5 and kids tickets cost $2. if 295 tickets wer sold and a total of $950 was collected how many students tickets were sold?

Answer by stanbon(75887)

write an equation for
adult tickets to a play cost $5 and kids tickets cost $2. if 295 tickets wer sold and a total of $950 was collected how many students tickets were sold?
-----------------
a + k = 295 tickets
5a+2k = 950 dollars
----------------------------
Substitute for "a" and solve for "k":
5(295-k)+2k = 950
1475 - 5k + 2k = 950
1475 - 3k = 950
3k = 525
k = 175 (# of kid tickets sold)
----
Since a + k = 295, a = 120 (# of adult tickets sold)
====================================================
Cheers,
Stan H.